# Look, Listen, and Answer: Overcoming Biases for Audio-Visual Question Answering

### v1

@@ Abstract
Audio-Visual Question Answering (AVQA) is a complex multi-modal reasoning task, demanding intelligent systems to accurately respond to natural language queries based on audio-video input pairs. Nevertheless, prevalent AVQA approaches are prone to overlearning dataset biases, resulting in poor robustness. Furthermore, current datasets may not provide a precise diagnostic for these methods. To tackle these challenges, firstly, we propose a novel dataset, MUSIC-AVQA-R, crafted in two steps: rephrasing questions within the test split of a public dataset (MUSIC-AVQA) and subsequently introducing distribution shifts to split questions. The former leads to a large, diverse test space, while the latter results in a comprehensive robustness evaluation on rare, frequent, and overall questions. Secondly, we propose a robust architecture that utilizes a multifaceted cycle collaborative debiasing strategy to overcome bias learning. Experimental results show that this architecture achieves state-of-the-art performance on both datasets, especially obtaining a significant improvement of 9.68% on the proposed dataset. Extensive ablation experiments are conducted on these two datasets to validate the effectiveness of the debiasing strategy. Additionally, we highlight the limited robustness of existing multi-modal QA methods through the evaluation on our dataset.
Look, Listen, and Answer: Overcoming
Biases for Audio-Visual Question Answering
Jie Ma1, Min Hu1,4, Pinghui Wang1, Wangchun Sun1, Lingyun Song2, Hongbin Pei1, Jun Liu3, Youtian Du1 1 MOE KLINNS Lab, Xi’an Jiaotong University, China 2 School of Computer Science, Northwestern Polytechnical University, China 3 School of Computer Science and Technology, Xi’an Jiaotong University, China 4 China Mobile System Integration Co.
@@ 1 Introduction
Humans possess the extraordinary capacity to seamlessly integrate auditory and visual cues, effectively establishing a cohesive relationship between visual and auditory stimuli Lin et al. (2023). Audio-Visual Question Answering (AVQA) Alamri et al. (2019); Yang et al. (2022b); Li et al. (2022a); Yun et al. (2021); Yang et al. (2022b) seeks to enable intelligent systems to acquire this capability and produce answers based on provided natural language questions. It requires the system to learn high-order interaction representations of the concepts encompassed with audio, video, and language modalities. As is known to us Wen et al. (2021), the high-level reasoning ability of the system mainly relies on large-scale data that does not contain harmful biases or statistical regularities.
@@
Figure 1: The question in current AVQA datasets is generated by a limited set of predefined templates, which is not in line with the real-world scenario. Moreover, the current dataset has not yet achieved refined evaluations, such as performance measurements on rare samples. Our findings indicate that existing methods such as STG Li et al. (2022a) are not robust in answering rephrased questions, which may be attributed to excessive bias learning, such as memorizing statistical regularities between critical question words and answers.
@@
Nevertheless, completely avoiding the negative bias in datasets seems challenging. Previous studies Li et al. (2023); Ravichander et al. (2023); Miller et al. (2020); Agrawal et al. (2018) in visual and extractive QA have investigated the bias from the perspective of changing answer distributions and human-in-the-loop adversarial attacks. Drawing inspiration from these works, several open questions are proposed for the AVQA task, concerning model evaluations and model designs.
@@
Question 1: have existing datasets comprehensively measured model robustness? The questions in the current AVQA dataset Schwartz et al. (2019); Yun et al. (2021) are generated by a limited set of predefined templates, such as the 33 templates in the MUSIC-AVQA dataset Li et al. (2022a). Fig. 1 shows the samples in the training and test split, which are produced using a predefined template. The only difference between the samples is a single word, which may not be consistent with the real-world situation. Furthermore, current datasets cannot reflect the performance on rare or less common samples, which is an important indicator for evaluating model robustness.
@@
Question 2: have existing methods overcome the data bias? We found that existing methods such as STG Li et al. (2022a) are brittle for the question with rare answers. This could be attributed to memorizing the statistical regularity between critical question words and answers, such as the connection between “Is”, “Playing”, and “Yes”.
@@
In this paper, we present the development of a novel dataset called MUSIC-AVQA-R, which aims to address the first question precisely. The dataset complements MUSIC-AVQA Li et al. (2022a) and provides a more refined diagnostic for current AVQA methods. To preserve the inherent bias, we maintain the original training and validation splits of the MUSIC-AVQA dataset. In contrast, we employ a human-machine collaboration mechanism to rephrase the question in the test split. This ensures diverse and natural question forms while remarkably expanding the number of questions from 9,129 to 211,572. We introduce a distribution shift based on answer distributions of specific question types. This allows us to measure performance on both frequent (in-distribution) and rare (out-of-distribution) data simultaneously.
@@
To tackle the second question, we propose a robust framework that applies a Multifaceted Cycle Collaborative Debiasing (MCCD) strategy. Specifically, the strategy introduces a novel optimization objective, which enlarges the distribution difference between uni-modal (questions, audio, and video) and multi-modal logit. By doing so, our model becomes less prone to learning biases from individual modalities effectively. Intuitively, we cannot choose the correct answer based on only one modality. Hence, MCCD employs cycle guidance to constrain the logit distribution of each modality, thereby promoting the similarity of uni-modal logit distribution. The experimental results demonstrate that our framework yields significant improvements on both datasets, with a particularly notable enhancement of 9.68% observed on the MUSIC-AVQA-R dataset.
In summary, our contributions are fourfold:
  • We propose a novel dataset MUSIC-AVQA-R and a set of respective evaluation metrics. This enables us to thoroughly evaluate the reasoning behavior of AVQA models and characterize their generalization capabilities in in- and out-of-distribution scenarios.
  • We present an AVQA architecture that incorporates the MCCD strategy to overcome training biases. To the best of our knowledge, this is the first work to systematically explore biases in the AVQA task from model evaluations as well as model designs.
  • We conduct extensive experiments on MUSIC-AVQA and MUSIC-AVQA-R to verify the effectiveness and superiority of our proposed architecture and debiasing strategy.
  • We evaluate 12 recent multimodal QA methods on the proposed dataset and show their limited ability to generalize not only in-distribution scenarios but also in out-of-distribution situations.
@@ 2.1 Model Robustness Evaluation
Despite the notable achievements of QA datasets Rajpurkar et al. (2016); Goyal et al. (2019); Hudson and Manning (2019); Yang et al. (2022a); Li et al. (2022b), they suffer from biases, resulting in incomplete evaluations. In recent years, numerous studies have tackled this issue from various perspectives Ma et al. (2023).
One avenue of research Agrawal et al. (2018); Ko et al. (2020); Dancette et al. (2021) reorganizes existing datasets, thereby making the distribution between training and testing splits significantly different or even reversed. The reorganized datasets reflect the performance in the out-of-distribution situation but lack measurement in the in-distribution scenario. To this end, GQA-OOD Kervadec et al. (2021) introduces the distribution shift in both the validation and test splits to assess visual QA models in in- and out-of-distribution scenarios simultaneously. Nevertheless, its question is generated using pre-defined templates, which may not accurately reflect real-world scenarios. In addition, the number of questions in the GQA-OOD test split is only 2,796, which may not reflect the real generalization ability of visual QA models due to the presence of a limited number of testing samples Harrell Jr et al. (1996).
Inspired by the adversarial attack, another line of works Sheng et al. (2021); Li et al. (2021b) regard the dataset construction as a game played by two parties: a human annotator and a well-trained model. Only samples generated by humans that successfully attack the model are incorporated into the dataset. Different from the mentioned works, our dataset, MUSIC-AVQA-R, not only prioritizes question diversity but also considers the volume of test samples. Notably, modifications are exclusively applied to the test split, as opposed to simultaneous adjustments in both validation and test splits. These enhance the precision and comprehensiveness of robustness evaluation. Furthermore, our dataset is the first AVQA dataset focusing on robustness evaluation to the best of our knowledge.
@@ 2.2 Bias Dependency Elimination
A variety of debiasing QA methods have been proposed to overcome bias memorization. These methods can be divided into four classes Ma et al. (2023): ensemble learning, data augmentation, contrastive learning, and answer re-ranking.
Ensemble learning methods Cadene et al. (2019); Ko et al. (2020); Niu and Zhang (2021); Wen et al. (2021); Cho et al. (2023) typically leverage a combination of a bias learner and a vanilla QA model to comprehensively predict answers. Data augmentation methods Kv and Mittal (2020); Abbasnejad et al. (2020); Chen et al. (2022); Teney et al. (2021) generate additional question-answer pairs to balance the data distribution. Based on the positive and negative sample generation, contrastive learning-based methods Liang et al. (2020); Zhu et al. (2021); Si et al. (2022) strive to learn an embedding space where similar sample pairs are closely clustered while disparate ones are distinctly separated. Consequently, the vanilla QA method is optimized jointly through contrastive and QA losses. Answer re-ranking methods Wu and Mooney (2019); Jing et al. (2020); Shrestha et al. (2020); Gat et al. (2020); Si et al. (2021) primarily focus on reordering the answers predicted by the vanilla QA model to enhance context comprehension, such as vision grounding.
To the best of our knowledge, COCA Lao et al. (2023) is the only work to mitigate the bias learning in the AVQA task, which first employs causal regularization to intervene bias-irrelevant causal effects and then introspects predictions. Unlike the mentioned works, which only consider language biases, our method considers audio, vision, language biases, and their collaboration.
In recent years, pre-trained multimodal models Lu et al. (2019); Wu and Mooney (2019); Tan and Bansal (2019); Li et al. (2021a); Lin et al. (2023); Girdhar et al. (2023); Zhang et al. (2023) have achieved notable progress in various tasks. However, recent studies He et al. (2020); Ma et al. (2023) found that they also suffer from biases. Our proposed architecture is compatible with these models, thereby improving their robustness.
@@
3 Dataset Creation and Analysis
We introduce the first dataset, MUSIC-AVQA-R, to evaluate the robustness of AVQA models. The construction of this dataset involves two key processes: rephrasing and splitting. The former involves the rephrasing of original questions generated from pre-defined question templates, and the latter is dedicated to the categorization of questions into frequent (head) and rare (tail) subsets.
@@ 3 Dataset Creation and Analysis
3.1 Rephrasing
The questions within the existing dataset Yun et al. (2021); Li et al. (2022a) are formulated using a restricted collection of pre-defined templates. To augment diversity and reality, we employ a rephrasing tool, such as ChatGPT, to rephrase each question within the test split of MUSIC-AVQA 25 times. To ensure the rephrasing quality, three volunteers participate in a verification process where their consensus through voting is required. Rephrasings are incorporated into the dataset only when two or more individuals validate the quality of the modifications. According to our statistics, 92.4% of rephrasings pass this validation. Fig. 2 illustrates the distribution of rephrased questions based on their initial three words. We see that our rephrasing questions have various formats, which are in line with real-world situations. Furthermore, an expansion of the question count within the test split has been implemented, escalating from 9,129 to 211,572 questions. This substantial increase in the volume of test samples enhances the precision of evaluations for AVQA methods.
3.2 Splitting
To provide a precise diagnostic for AVQA models, we introduce a distribution shift based on answer distributions of specific question types, following Kervadec et al. (2021). Guided by this distribution, we categorize rephrased questions into head and tail, enabling the assessment of in-distribution and out-of-distribution performance, respectively. We also utilize the overall performance to assess the model effectiveness on the entire test split.
Finally, we categorize the samples into head and tail classes. We define the tail class as class $i$ with $|a_{i}|\leq 1.2\mu(a)$ following Kervadec et al. (2021), where $|a_{i}|$ represents the number of samples belonging to answer class $i$, and $\mu(a)$ denotes the average sample count for a group. Consequently, the tail samples are rare, while the head samples are more prevalent within a group. Fig. 3(b) illustrates the statistics of head and tail samples across various groups within each task.
The aforementioned procedure is only carried out in the test split of MUSIC-AVQA. Consequently, our proposed dataset, MUSIC-AVQA-R, allows for a more precise and comprehensive evaluation of models handling data biases.
@@ 4 Method
To mitigate bias learning, we propose a robust AVQA architecture that integrates a multifaceted cycle collaborative debiasing strategy. Fig. 4 illustrates an overview of our proposed architecture. It first learns the uni-modal and multimodal representations using a pre-trained model. Then, the architecture utilizes distinct bias learners to capture uni-modal biases. Finally, a collaborative debiasing strategy is leveraged to magnify the disparity between fusion logit and bias logit, obtained based on multi-modality and uni-modality representations, respectively. Meanwhile, a cycle guidance mechanism is employed to maintain the similarity between bias logit.
Uni-modal Embedding. Given an AVQA sample comprising a video sequence and a corresponding question, we initially partition the sequence, consisting of visual and audio tracks, into $T$ non-overlapping pairs of visual and audio segments, denoted as $\{V_{t},A_{t}\}_{t=1}^{T}$, where each segment spans one second. Subsequently, a distinct embedding layer is employed to acquire uni-modal embeddings. Specifically, we employ a pre-trained VGGish model Gemmeke et al. (2017) with fixed parameters, which is a VGG-like audio processing network, to obtain an audio embedding vector. For video embedding vectors, we employ a pre-trained ResNet-18 with fixed parameters on the frames. The VisualBert model Li et al. (2019a) is applied to obtain a word-level question embedding vector. To ensure dimension matching, distinct linear layers are applied to the aforementioned vectors, resulting in uni-modal embeddings $\mathbf{A}_{i}^{\mathrm{e}},\mathbf{V}_{i}^{\mathrm{e}}\in\mathbb{R}^{T\times 768}$ and $\mathbf{Q}_{i}^{\mathrm{e}}\in\mathbb{R}^{l\times 768}$, where $l$ is the question length.
@@
$\mathcal{L}_{\mathrm{qa}}$ $=\frac{1}{K}\sum\hat{\mathbf{y}}_{i}^{\mathrm{q}}\left(\log\hat{\mathbf{y}}_{i}^{\mathrm{q}}-\log\hat{\mathbf{y}}_{i}^{\mathrm{a}}\right),$ (3)
where $\beta$ is the factor to control weight, $\mathcal{L}_{\mathrm{qa}}$ denotes the relative entropy between the question and audio logit, and $\mathcal{L}_{\mathrm{av}}$ and $\mathcal{L}_{\mathrm{vq}}$ are obtained similarly.
Finally, we utilize the summation of $\mathcal{L}_{\mathrm{d}}$, $\mathcal{L}_{\mathrm{c}}$ and $\mathcal{L}_{\mathrm{a}}$ to optimize the parameters of our method, where $\mathcal{L}_{\mathrm{a}}$ is the loss of answer prediction that is regarded as a multi-classification problem. The training details are shown in Appendix A.

### v3

@@ Abstract
Audio-Visual Question Answering (AVQA) is a complex multi-modal reasoning task, demanding intelligent systems to accurately respond to natural language queries based on audio-video input pairs. Nevertheless, prevalent AVQA approaches are prone to overlearning dataset biases, resulting in poor robustness. Furthermore, current datasets may not provide a precise diagnostic for these methods. To tackle these challenges, firstly, we propose a novel dataset, MUSIC-AVQA-R, crafted in two steps: rephrasing questions within the test split of a public dataset (MUSIC-AVQA) and subsequently introducing distribution shifts to split questions. The former leads to a large, diverse test space, while the latter results in a comprehensive robustness evaluation on rare, frequent, and overall questions. Secondly, we propose a robust architecture that utilizes a multifaceted cycle collaborative debiasing strategy to overcome bias learning. Experimental results show that this architecture achieves state-of-the-art performance on MUSIC-AVQA-R, notably obtaining a significant improvement of 9.32%. Extensive ablation experiments are conducted on the two datasets mentioned to analyze the component effectiveness within the debiasing strategy. Additionally, we highlight the limited robustness of existing multi-modal QA methods through the evaluation on our dataset. We also conduct experiments combining various baselines with our proposed strategy on two datasets to verify its plug-and-play capability. Our dataset and code are available at https://github.com/reml-group/MUSIC-AVQA-R.
@@ 1 Introduction
Humans possess the extraordinary capacity to seamlessly integrate auditory and visual cues, effectively establishing a cohesive relationship between visual and auditory stimuli [1]. Audio-Visual Question Answering (AVQA) [2, 3, 4, 5, 3] seeks to enable intelligent systems to acquire this capability and produce answers based on provided natural language questions. It requires the system to learn high-order interaction representations of the concepts encompassed with audio, video, and language modalities. As is known to us [6, 7, 8], the high-level reasoning ability of the system mainly relies on large-scale data that does not contain harmful biases or statistical regularities.
@@
Figure 1: The question in current AVQA datasets is generated by a limited set of predefined templates, which may not be in line with the real-world scenario. Our findings indicate that existing methods [5, 1] such as STG [4] are not robust, which may be attributed to excessive bias learning, such as memorizing statistical regularities between critical question words and answers.
@@
Nevertheless, completely avoiding the negative bias in datasets seems challenging. Previous studies [9, 10, 11, 12] in visual and extractive QA have investigated the bias from the perspective of changing answer distributions and human-in-the-loop adversarial attacks. Drawing inspiration from these works, several open questions are proposed for the AVQA task, concerning model evaluations and model designs.
@@
Question 1: have existing datasets comprehensively measured model robustness? The questions in the current AVQA dataset [13, 5, 14, 3] are generated by a limited set of predefined templates, such as the 33 templates in the MUSIC-AVQA dataset [4]. Fig. 1 shows the samples in the training and test split, which are produced using a predefined template. The observed difference mainly stems from a single word, leading to a limited vocabulary size of only 93 words. This has the potential to deviate from real-world scenarios. Moreover, current datasets cannot reflect the performance on rare or less common samples, which is an important indicator for evaluating model robustness [15, 16].
@@
Question 2: have existing methods overcome the data bias? We found that existing methods [5, 1, 17, 18] such as STG [4] are brittle for the question with rare answers. This may be attributed to memorizing the statistical regularity between critical question words and answers, such as the connection between “Is”, “Playing”, and “Yes”. Specifically, the experimental result [4] shows that STG achieves an accuracy of 54.09% on the test split of MUSIC-AVQA only given questions.
@@
In this paper, we present the development of a novel dataset called MUSIC-AVQA-R, which aims to address the first question precisely. The dataset complements MUSIC-AVQA [4] and provides a more refined diagnostic for current AVQA methods. To preserve the inherent bias, we maintain the original training and validation splits of the MUSIC-AVQA dataset. In contrast, we employ a human-machine collaboration mechanism to rephrase the question in the test split. This ensures diverse and natural question forms while remarkably expanding the number of questions from 9,129 to 211,572. We introduce a distribution shift based on answer distributions of specific question types. This allows us to measure performance on both frequent (in-distribution) and rare (out-of-distribution) data simultaneously.
@@
To tackle the second question, we propose a robust framework that applies a Multifaceted Cycle Collaborative Debiasing (MCCD) strategy. Specifically, the strategy introduces a novel optimization objective, which enlarges the distribution difference between uni-modal (question, audio, and video) and multi-modal logit. By doing so, our model becomes less prone to learning biases from individual modalities. Intuitively, we cannot choose the correct answer based on only one modality. Hence, MCCD employs cycle guidance to constrain the logit distribution of each modality, thereby promoting the similarity of uni-modal logit distribution. The experimental results demonstrate that our framework yields significant improvements on both datasets, with a particularly notable enhancement of 9.32% observed on the MUSIC-AVQA-R dataset.
In summary, our contributions are fourfold: (1) We propose a novel dataset MUSIC-AVQA-R and a set of respective evaluation metrics. This enables us to thoroughly evaluate the reasoning behavior of AVQA models and characterize their generalization capabilities in in- and out-of-distribution scenarios. (2) We present an AVQA architecture that incorporates the MCCD strategy to overcome training biases. To the best of our knowledge, this is the first work to systematically explore biases in the AVQA task from model evaluations as well as model designs. (3) We conduct extensive experiments on MUSIC-AVQA and MUSIC-AVQA-R to verify the effectiveness and superiority of our proposed architecture and debiasing strategy. (4) We evaluate 13 recent multimodal QA methods on the proposed dataset and show their limited ability to generalize not only in-distribution scenarios but also in out-of-distribution situations.
@@ 2.1 Model Robustness Evaluation
Despite the notable achievements of QA datasets [19, 20, 21, 22, 23, 3], they suffer from biases, resulting in incomplete evaluations. In recent years, numerous studies have tackled this issue from various perspectives [24, 25, 26, 27].
One avenue of research [12, 28, 29] reorganizes existing datasets, thereby making the distribution between training and testing splits significantly different or even reversed. The reorganized datasets reflect the performance in the out-of-distribution situation but lack measurement in the in-distribution scenario. To this end, GQA-OOD [30] introduces the distribution shift in both the validation and test splits to assess visual QA models in both scenarios simultaneously. Nevertheless, the number of questions in the GQA-OOD test split is only 2,796, which may not reflect the real generalization ability of visual QA models due to the presence of a limited number of testing samples [31]. Inspired by the adversarial attack, another line of works [32, 33] regard the dataset construction as a game played by two parties: a human annotator and a well-trained model. Only samples generated by humans that successfully attack the model are incorporated into the dataset. In addition, there exists another line of work [14] that complements videos and questions to obtain balanced training data.
Different from the mentioned works, our dataset, MUSIC-AVQA-R, not only prioritizes question diversity but also considers the volume of test samples. This enhances the precision and comprehensiveness of robustness evaluation. Moreover, we recognize the formidable challenge of obtaining completely pure training data. As such, we opt to retain the inherent bias present in both the training and validation splits. Our primary objective is to inspire the community to enhance model robustness through the implementation of debiasing strategies, rather than striving for balanced training data. Remarkably, to the best of our knowledge, our dataset is the first AVQA dataset explicitly designed for robustness evaluation.
@@ 2.2 Bias Dependency Elimination
A variety of debiasing QA methods [34, 35, 36, 37] have been proposed to overcome bias memorization. These methods can be divided into four classes [24]: ensemble learning, data augmentation, contrastive learning, and answer re-ranking.
Ensemble learning methods [38, 28, 39, 6, 40, 41] typically leverage a combination of a bias learner and a vanilla QA model to comprehensively predict answers. Data augmentation methods [42, 43, 44, 45] generate additional question-answer pairs to balance the data distribution. Based on the positive and negative sample generation, contrastive learning-based methods [46, 47, 48] strive to learn an embedding space where similar sample pairs are closely clustered while disparate ones are distinctly separated. Consequently, the vanilla QA method is optimized jointly through contrastive and QA losses. Answer re-ranking methods [34, 49, 50, 51, 52] primarily focus on reordering the answers predicted by the vanilla QA model to enhance context comprehension, such as vision grounding.
To the best of our knowledge, COCA [53] is the only work to mitigate the bias learning in the AVQA task, which first employs causal regularization to intervene bias-irrelevant causal effects and then introspects predictions. Unlike the mentioned works, which only consider language biases, our method considers audio, vision, language biases, and their collaboration. The proposed MCCD strategy features plug-and-play capability, enhancing the debiasing potential of baseline methods.
3 Dataset Creation and Analysis
We introduce the first dataset, MUSIC-AVQA-R, to evaluate the robustness of AVQA models. The construction of this dataset involves two key processes: rephrasing and splitting. The former involves the rephrasing of questions in the test split of MUSIC-AVQA, and the latter is dedicated to the categorization of questions into frequent (head) and rare (tail) subsets.
3.1 Rephrasing
The questions within the existing dataset [5, 4] are formulated using a restricted collection of pre-defined templates. To augment diversity and reality, we employ a rephrasing tool [^1] https://quillbot.com/paraphrasing-tool to rephrase each question 25 times. To ensure the rephrasing quality, three annotators participate in a verification process where their consensus through voting is required. They are all senior students in the field of information science, with one specializing in computer science and the other two in automation. Their extensive professional background equips them with the ability to assess whether the above rephrasing fulfills the requirement. Rephrasings are incorporated into the dataset only when two or more individuals validate the quality of the modifications. According to the statistics, 92.4% of rephrasings pass this validation, and the Fleiss Kappa value used to measure vote consistency is 0.839. Please see details in Table 4 of Appendix B. These results strongly suggest an exceptionally high quality of the rephrasing efforts. Fig. 2 illustrates the distribution of rephrased questions based on their initial three words. We see that our rephrasing questions have various formats, and the comparison between the two datasets is shown in Fig. 6 of Appendix B. The vocabulary size of our dataset is 465, which is 5x larger than MUSIC-AVQA. These results indicate that our dataset is more in line with the real-world scenario. Furthermore, an expansion of the question count within the test split has been implemented, escalating from 9,129 to 211,572 questions. This substantial increase in the volume of test samples enhances the precision of evaluations for AVQA methods.
@@ 3.1 Rephrasing
3.2 Splitting
To provide a precise diagnostic for AVQA models, we introduce a distribution shift based on answer distributions of specific question types, following [30]. Guided by this distribution, we categorize rephrased questions into head and tail, enabling the assessment of in-distribution and out-of-distribution performance, respectively. We also utilize the overall performance to assess the model effectiveness on the entire test split.
@@ 3.2 Splitting
Finally, we categorize the samples into head and tail classes. We define the tail class as class $i$ with $|a_{i}|\leq 1.2\mu(a)$ following [30], where $|a_{i}|$ represents the number of samples belonging to answer class $i$, and $\mu(a)$ denotes the average sample count for a group. Consequently, the tail samples are rare, while the head samples are more prevalent within a group. Fig. 3(b) illustrates the statistics of head and tail samples across various groups within each task.
@@ 4 Method
To mitigate bias learning, we propose a robust AVQA architecture that integrates a multifaceted cycle collaborative debiasing strategy. Fig. 4 illustrates an overview of our proposed architecture. It first learns the uni-modal and multimodal representations using a pre-trained model. Then, the architecture utilizes distinct bias learners to capture uni-modal biases. Finally, a collaborative debiasing strategy is leveraged to magnify the disparity between fusion logit and bias logit, obtained based on multi-modality and uni-modality representations, respectively. Meanwhile, a cycle guidance mechanism is employed to maintain the similarity between bias logit. The aforementioned procedure is only carried out in the test split of MUSIC-AVQA. Consequently, our proposed dataset, MUSIC-AVQA-R, allows for a more precise and comprehensive evaluation of models handling data biases.
Uni-modal Embedding. Given an AVQA sample comprising a video sequence and a corresponding question, we initially partition the sequence, consisting of visual and audio tracks, into $T$ non-overlapping pairs of visual and audio segments, denoted as $\{V_{t},A_{t}\}_{t=1}^{T}$, where each segment spans one second. Subsequently, a distinct embedding layer is employed to acquire uni-modal embeddings. Specifically, we employ a pre-trained VGGish model [54] with fixed parameters, which is a VGG-like audio processing network, to obtain an audio embedding vector. For video embedding vectors, we employ a pre-trained ResNet-18 with fixed parameters on the frames. The VisualBert model [55] is applied to obtain a word-level question embedding vector. To ensure dimension matching, distinct linear layers are applied to the aforementioned vectors, resulting in uni-modal embeddings $\mathbf{A}_{i}^{\mathrm{e}},\mathbf{V}_{i}^{\mathrm{e}}\in\mathbb{R}^{T\times 768}$ and $\mathbf{Q}_{i}^{\mathrm{e}}\in\mathbb{R}^{l\times 768}$, where $l$ is the question length.
@@
where $\beta$ is the factor to control weight, $\mathcal{L}_{\mathrm{qa}}=\frac{1}{K}\sum_{i=1}^{K}\hat{\mathbf{y}}_{i}^{\mathrm{q}}\left(\log\hat{\mathbf{y}}_{i}^{\mathrm{q}}-\log\hat{\mathbf{y}}_{i}^{\mathrm{a}}\right)$ denotes the relative entropy between the question $\hat{\mathbf{y}}_{i}^{\mathrm{q}}$ and audio logit $\hat{\mathbf{y}}_{i}^{\mathrm{a}}$, $\mathcal{L}_{\mathrm{av}}=\frac{1}{K}\sum_{i=1}^{K}\hat{\mathbf{y}}_{i}^{\mathrm{a}}\left(\log\hat{\mathbf{y}}_{i}^{\mathrm{a}}-\log\hat{\mathbf{y}}_{i}^{\mathrm{v}}\right)$ is the relative entropy between the audio $\hat{\mathbf{y}}_{i}^{\mathrm{a}}$ and video logit $\hat{\mathbf{y}}_{i}^{\mathrm{v}}$, and $\mathcal{L}_{\mathrm{vq}}=\frac{1}{K}\sum_{i=1}^{K}\hat{\mathbf{y}}_{i}^{\mathrm{v}}\left(\log\hat{\mathbf{y}}_{i}^{\mathrm{v}}-\log\hat{\mathbf{y}}_{i}^{\mathrm{q}}\right)$ represents the relative entropy between the video $\hat{\mathbf{y}}_{i}^{\mathrm{v}}$ and question logit $\hat{\mathbf{y}}_{i}^{\mathrm{q}}$.
Finally, we utilize the summation of $\mathcal{L}_{\mathrm{d}}$, $\mathcal{L}_{\mathrm{c}}$ and $\mathcal{L}_{\mathrm{a}}$ to optimize the parameters of our method. $\mathcal{L}_{\mathrm{a}}=-\frac{1}{K}\sum_{i=1}^{K}\mathbf{y}_{i}^{\mathrm{f}}\log\hat{\mathbf{y}}_{i}^{\mathrm{f}}$ is the loss of answer prediction that is regarded as a multi-classification problem, where $\mathbf{y}_{i}^{\mathrm{f}},\hat{\mathbf{y}}_{i}^{\mathrm{f}}$ denote the one-hot answer label and logit of multi-modality fusion, respectively. The training details are shown in Appendix A.